DQuestion 1 2 pts Select the statements below that are TRUE. Kc and Kp are the same thing as the k (rate constant) that you learned about before. Kc has units of molarity and Kp has units of pressure D Kp uses the same equation as Kc, but with pressure instead of concentration Kc is equal to Kp when the number of moles of gas (n) doesnt change. Solids and liquid are not included in the Kc and Kp equations because they have an activity equal to 1. If there is a liquid in a reaction, you use the vapor press for that entry in the Kp equation.
Answer #1

g 1 @ kp = kc (RT) an kp and kc are not same. honce False → K - cinit as Conco as pressur kp = unit A + B c * = [c] [A] [B] kALG) - CEg + 2 B 1g) 2 I 2 - - 22 + 2 C ta 2 -X 2-22 At equilibrium (P = 0.5 atm c = 0.5 atm PA – 2- - 2-0.5 IPA = 115 atm PB